OpenStudy (anonymous):

MEDAL!! There are 20 players on a hockey team. 5 of them are rookies. Determine the probability that none of the rookies will be randomly selected to be included in the team's group of 6 starters. A. 15/114 B. 1/3,876 C. 1/38,760 D. 1001/7752

OpenStudy (anonymous):

would my answer be A?

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

\[\frac{ \left(\begin{matrix}5 \\ 0\end{matrix}\right)\left(\begin{matrix}15 \\ 5\end{matrix}\right) }{ \left(\begin{matrix}20 \\ 5\end{matrix}\right) }\]

OpenStudy (anonymous):

this is because you really have two groups... rookies and non-rookies. there are 5 rookies and 15 non-rookies. you are sampling witthout replacement and order doesn't matter so you use combinations. for the denominator, there are 20 players, and you need to choose 5. for the numerator, there are 5 rookies, of which you choose 0 and there are 15 non-rookies, of which you choose 5. I hope that helps!

OpenStudy (anonymous):

yea:) so my answer would be A correct?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

.19369.... which is not A

OpenStudy (anonymous):

hmmmm, should I solve that fraction you gave me, and will that give me my answer?

OpenStudy (anonymous):

oops, hold on I may have made a calc error...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it =21699 which equals 0.44017 but cant figure out howw to put it into a fraction

OpenStudy (anonymous):

i found my mistake... there are 6 player starting, not 5. just a sec...

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

=.129127... = > still not A I suggest you compute the combinations. do you have a calculator to do that?

OpenStudy (anonymous):

oe sec

OpenStudy (anonymous):

15/20 *14/19 *13/18 *12/17 *11/16 *10/15 = 1001/7752

OpenStudy (anonymous):

Yeah, that's another way to do it! Great for you!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you're so welcome!